# Accelerated Test-Time Scaling with Model-Free Speculative Sampling

### v1

@@ 5.1 Main results
Table 1: Speculative decoding performance in multi-trajectory reasoning. We report the average throughput (T) and acceptance length (A) for multi-trajectory test-time scaling scenarios, with different number of reasoning trajectories per problem. We evaluate each model on AIME-2024 (AIME), GPQA-Diamond (GPQA), and LiveCodeBench (LCB). The best values are highlighted in bold.
4 Trajectories 8 Trajectories 16 Trajectories
AIME GPQA LCB AIME GPQA LCB AIME GPQA LCB
DeepSeek-R1-Distill-Qwen-7B
Plain T 26.63 31.34 27.75 26.63 31.34 27.75 26.63 31.34 27.75
Eagle-2 T 29.91 31.69 27.61 29.91 31.69 27.61 29.91 31.69 27.61
A 2.21 1.99 2.13 2.21 1.99 2.13 2.21 1.99 2.13
PLD T 43.93 50.49 44.01 44.95 53.04 45.08 46.60 53.47 46.02
A 1.78 1.81 1.73 1.84 1.89 1.79 1.89 1.96 1.85
ANPD T 45.52 57.39 46.30 46.40 58.97 47.86 47.06 60.25 48.81
A 1.89 1.97 1.88 1.92 2.03 1.91 1.96 2.11 1.96
SAM T 44.35 53.21 45.63 45.64 55.47 47.24 47.64 57.53 48.92
A 1.81 1.87 1.85 1.89 1.96 1.89 1.97 2.03 1.95
Recycle T 61.38 71.51 60.62 61.70 71.55 60.93 60.86 71.23 61.36
A 2.76 2.73 2.73 2.77 2.73 2.73 2.77 2.73 2.74
SAM + Recycle T 61.11 70.43 62.20 60.66 69.98 63.41 60.63 69.85 63.39
A 2.71 2.73 2.68 2.69 2.74 2.69 2.68 2.71 2.67
STAND (Ours) T 64.99 83.47 69.70 66.88 87.02 71.83 69.15 91.17 74.14
A 3.21 3.48 3.30 3.35 3.70 3.47 3.46 3.90 3.64
DeepSeek-R1-Distill-Qwen-14B
Plain T 17.76 18.16 17.43 17.76 18.16 17.43 17.76 18.16 17.43
Eagle-2 T 25.38 24.86 21.89 25.38 24.86 21.89 25.38 24.86 21.89
A 2.72 2.44 2.51 2.72 2.44 2.51 2.72 2.44 2.51
PLD T 24.37 26.60 23.36 25.44 27.36 23.96 26.35 28.43 24.97
A 1.74 1.82 1.74 1.84 1.91 1.81 1.92 2.00 1.88
ANPD T 25.74 28.21 24.78 26.12 29.51 25.63 26.49 30.62 26.32
A 1.87 1.97 1.87 1.91 2.04 1.93 1.96 2.13 1.99
SAM T 25.22 28.03 24.41 26.11 29.39 25.37 27.25 30.59 26.67
A 1.78 1.85 1.79 1.88 1.95 1.87 1.98 2.06 1.96
Recycle T 34.97 38.99 34.05 35.06 38.89 33.98 35.31 38.81 33.96
A 2.78 2.73 2.72 2.77 2.73 2.72 2.77 2.74 2.72
SAM + Recycle T 34.81 38.24 34.15 35.16 38.57 34.19 35.53 38.99 34.31
A 2.70 2.71 2.65 2.71 2.71 2.66 2.72 2.71 2.65
STAND (Ours) T 37.56 43.71 38.71 39.13 46.81 40.45 40.76 49.11 42.72
A 3.16 3.42 3.29 3.28 3.63 3.47 3.42 3.86 3.65
@@ 4 Trajectories 8 Trajectories 16 Trajectories
In Figure 1 and Table 1, we evaluate STAND’s performance in multi-trajectory inference, where we generate multiple candidate answers by sequentially producing k independent reasoning traces and then aggregate the results.
@@
As shown in Figure 1, STAND significantly improves decoding efficiency, achieving equivalent performance to plain decoding in less than 40% the time. Table 1 provides detailed throughput and acceptance length comparisons across methods. STAND not only achieves the highest throughput but also maintains longer acceptance lengths compared to baselines. Importantly, both metrics improve as we increase the number of trajectories, making STAND’s speedup advantage more pronounced with increased compute scaling.
@@ 5.2 Evaluation on single-trajectory decoding
Table 2: Single-trajectory evaluations. We report the throughput (T) and acceptance length (A) for generating a single sequence with DeepSeek-R1-Distill-Qwen-7B and 14B. The best values are highlighted in bold, and the runner-up is underlined.
AIME GPQA LCB T A T A T A DeepSeek-R1-Distill-Qwen-7B Plain 26.63 x 31.34 x 27.75 x Eagle-2 29.91 2.21 31.69 1.99 27.61 2.13 PLD 44.34 1.72 42.84 1.64 43.40 1.59 ANPD 46.18 1.88 54.05 1.82 44.79 1.80 SAM 40.85 1.69 48.45 1.69 42.92 1.80 Recycle 60.61 2.73 71.00 2.71 60.12 2.73 SAM + Recycle 61.15 2.70 71.51 2.81 62.78 2.69 Ours 61.79 3.07 75.39 3.05 66.41 3.01 DeepSeek-R1-Distill-Qwen-14B Plain 17.76 x 18.16 x 17.43 x Eagle-2 25.38 2.72 24.86 2.44 21.89 2.51 PLD 21.82 1.61 24.97 1.64 21.76 1.58 ANPD 25.60 1.76 26.40 1.79 23.16 1.76 SAM 23.26 1.63 25.38 1.65 22.36 1.63 Recycle 33.71 2.77 38.91 2.73 33.85 2.71 SAM + Recycle 34.35 2.67 37.53 2.72 34.45 2.70 Ours 34.52 2.91 38.71 3.00 34.86 2.93
While STAND is primarily designed to leverage information across multiple reasoning trajectories, we also evaluate its performance on single-trajectory generation, where the model only produces one long reasoning chain. As shown in Table 2, STAND achieves both the highest acceptance length and throughput in most scenarios, demonstrating its effectiveness even when generating individual solutions.
@@ 5.3 Ablations and analysis
Table 3: Effect of Stochastic Drafting. We report the throughput for generating 4 sequences with DeepSeek-R1-Distill-Qwen-7B, on AIME-2024.
AIME GPQA LCB T A T A T A Deterministic 62.13 2.94 73.67 2.98 63.44 2.90 Stochastic 63.44 3.24 81.20 3.56 65.90 3.29 + Gumbel-Top-K 64.99 3.21 83.47 3.48 69.70 3.30
@@
In Table 3, we compare three drafting approaches: deterministic drafting, basic stochastic drafting (using PyTorch’s multinomial sampling), and our optimized stochastic drafting with Gumbel-Top-K. For fair comparison, we separately perform tree optimization for determinisic drafting and stochastic drafting. Stochastic drafting consistently achieves higher acceptance lengths across all tasks, resulting in improved throughput compared to deterministic drafting. Our Gumbel-Top-K optimization further improves performance by maintaining similar acceptance lengths while significantly reducing latency, leading to even higher throughput.
Table 4: Effect of tree optimization. Comparison of mean acceptance lengths when generating 4 sequences with DeepSeek-R1-Distill-Qwen-7B on two datasets: AIME-2024 and GPQA-Diamond. We compare two types of static trees: the heuristic trees from Token Recycle and our data-optimized trees (optimized on AIME-2024). AIME-2024 represents in-domain (IND) performance since it was used for tree optimization, while GPQA-Diamond tests out-of-domain (OOD) generalization. The best values are highlighted in bold.
AIME (IND) GPQA (OOD) Heuristic Optimized Heuristic Optimized Throughput 59.96 64.99 77.32 83.47 Acc. Lens 3.17 3.21 3.35 3.48
@@
In Table 4, we showcase the effectiveness of our tree optimization technique. Specifically, we compare the performance of a heuristic tree originally used by Token Recycle (Luo et al., 2024) with our tree, optimized on the AIME-2024 dataset. The results demonstrates that the optimized tree improves performance on both AIME-2024 and GPQA-Diamond, showcasing that the optimization not only works for the dataset in the same domain, but also generatlizes to out-of-domain (OOD) tasks.
@@
We analyze how different drafting approaches lead to different optimal tree structures by comparing trees optimized for STAND versus Token Recycle. As shown in Figure 5, the tree optimized for STAND reaches greater depths, extending to 13 levels compared to 7 levels in the Token Recycle-optimized tree. This difference likely stems from STAND’s higher acceptance rate, which favors deeper, narrower tree structures under the same tree size budget.
@@ 6 Conclusion
In this work, we introduced STAND, a model-free speculative decoding approach that accelerates language model reasoning while maintaining accuracy. By utilizing reasoning trajectory redundancy and historical logit information, STAND significantly improves throughput over standard auto-regressive decoding. Our method outperforms existing alternatives, offering an efficient solution for scaling AI reasoning systems.
@@
While STAND demonstrates superior performance, it could be tested on more diverse reasoning patterns. Currently, the approach has been evaluated primarily on multi-sampling and single-trajectory reasoning scenarios. Although we expect similar benefits would extend to other test-time scaling methods like tree search and iterative refinement, it has not been tested yet.
Appendix A Appendix
A.1 Shape of the initial tree for optimization

### v2

@@ 5.1 Evaluation on multi-trajectory decoding
Table 1: Speculative decoding performance in multi-trajectory reasoning. We report the average throughput (T) and acceptance length (A) for multi-trajectory test-time scaling scenarios, with different number of reasoning trajectories per problem. We evaluate each model on AIME-2024 (AIME), GPQA-Diamond (GPQA), and LiveCodeBench (LCB). Best results are shown in bold.
4 Trajectories 8 Trajectories 16 Trajectories AIME GPQA LCB Avg. AIME GPQA LCB Avg. AIME GPQA LCB Avg. DeepSeek-R1-Distill-Qwen-7B Plain T 26.63 31.34 27.75 28.57 26.63 31.34 27.75 28.57 26.63 31.34 27.75 28.57 Eagle-2 T 29.91 31.69 27.61 29.74 29.91 31.69 27.61 29.74 29.91 31.69 27.61 29.74 (x1.12) (x1.01) (x0.99) (x1.04) (x1.12) (x1.01) (x0.99) (x1.04) (x1.12) (x1.01) (x0.99) (x1.04) A 2.21 1.99 2.13 2.11 2.21 1.99 2.13 2.11 2.21 1.99 2.13 2.11 PLD T 43.93 50.49 44.01 46.14 44.95 53.04 45.08 47.69 46.60 53.47 46.02 48.70 (x1.65) (x1.61) (x1.59) (x1.61) (x1.69) (x1.69) (x1.62) (x1.67) (x1.75) (x1.71) (x1.66) (x1.70) A 1.78 1.81 1.73 1.77 1.84 1.89 1.79 1.84 1.89 1.96 1.85 1.90 ANPD T 45.52 57.39 46.30 49.74 46.40 58.97 47.86 51.08 47.06 60.25 48.81 52.04 (x1.71) (x1.83) (x1.67) (x1.74) (x1.74) (x1.88) (x1.72) (x1.79) (x1.77) (x1.92) (x1.76) (x1.82) A 1.89 1.97 1.88 1.91 1.92 2.03 1.91 1.95 1.96 2.11 1.96 2.01 SAM T 44.35 53.21 45.63 47.73 45.64 55.47 47.24 49.45 47.64 57.53 48.92 51.36 (x1.67) (x1.70) (x1.64) (x1.67) (x1.71) (x1.77) (x1.70) (x1.73) (x1.79) (x1.84) (x1.76) (x1.80) A 1.81 1.87 1.85 1.84 1.89 1.96 1.89 1.91 1.97 2.03 1.95 1.98 Recycle T 61.38 71.51 60.62 64.50 61.70 71.55 60.93 64.73 60.86 71.23 61.36 64.48 (x2.30) (x2.28) (x2.18) (x2.26) (x2.32) (x2.28) (x2.20) (x2.27) (x2.29) (x2.27) (x2.21) (x2.26) A 2.76 2.73 2.73 2.74 2.77 2.73 2.73 2.74 2.77 2.73 2.74 2.75 SAM+Recycle T 61.11 70.43 62.20 64.58 60.66 69.98 63.41 64.68 60.63 69.85 63.39 64.62 (x2.29) (x2.25) (x2.24) (x2.26) (x2.28) (x2.23) (x2.29) (x2.26) (x2.28) (x2.23) (x2.28) (x2.26) A 2.71 2.73 2.68 2.71 2.69 2.74 2.69 2.71 2.68 2.71 2.67 2.69 STAND (Ours) T 64.99 83.47 69.70 72.72 66.88 87.02 71.83 75.24 69.15 91.17 74.14 78.15 (x2.44) (x2.66) (x2.51) (x2.55) (x2.51) (x2.78) (x2.59) (x2.63) (x2.60) (x2.91) (x2.67) (x2.74) A 3.21 3.48 3.30 3.33 3.35 3.70 3.47 3.51 3.46 3.90 3.64 3.67 DeepSeek-R1-Distill-Qwen-14B Plain T 17.76 18.16 17.43 17.78 17.76 18.16 17.43 17.78 17.76 18.16 17.43 17.78 Eagle-2 T 25.38 24.86 21.89 24.04 25.38 24.86 21.89 24.04 25.38 24.86 21.89 24.04 (x1.43) (x1.37) (x1.26) (x1.35) (x1.43) (x1.37) (x1.26) (x1.35) (x1.43) (x1.37) (x1.26) (x1.35) A 2.72 2.44 2.51 2.56 2.72 2.44 2.51 2.56 2.72 2.44 2.51 2.56 PLD T 24.37 26.6 23.36 24.78 25.44 27.36 23.96 25.59 26.35 28.43 24.97 26.58 (x1.37) (x1.46) (x1.34) (x1.39) (x1.43) (x1.51) (x1.37) (x1.44) (x1.48) (x1.57) (x1.43) (x1.49) A 1.74 1.82 1.74 1.77 1.84 1.91 1.81 1.85 1.92 2.00 1.88 1.93 ANPD T 25.74 28.21 24.78 26.24 26.12 29.51 25.63 27.09 26.49 30.62 26.32 27.81 (x1.45) (x1.55) (x1.42) (x1.48) (x1.47) (x1.63) (x1.47) (x1.52) (x1.49) (x1.69) (x1.51) (x1.56) A 1.87 1.97 1.87 1.90 1.91 2.04 1.93 1.96 1.96 2.13 1.99 2.03 SAM T 25.22 28.03 24.41 25.89 26.11 29.39 25.37 26.96 27.25 30.59 26.67 28.17 (x1.42) (x1.54) (x1.40) (x1.46) (x1.47) (x1.62) (x1.46) (x1.52) (x1.53) (x1.68) (x1.53) (x1.58) A 1.78 1.85 1.79 1.81 1.88 1.95 1.87 1.90 1.98 2.06 1.96 2.00 Recycle T 34.97 38.99 34.05 36.00 35.06 38.89 33.98 35.98 35.31 38.81 33.96 36.03 (x1.97) (x2.15) (x1.95) (x2.02) (x1.97) (x2.14) (x1.95) (x2.02) (x1.99) (x2.14) (x1.95) (x2.03) A 2.78 2.73 2.72 2.74 2.77 2.73 2.72 2.74 2.77 2.74 2.72 2.74 SAM+Recycle T 34.81 38.24 34.15 35.73 35.16 38.57 34.19 35.97 35.53 38.99 34.31 36.28 (x1.96) (x2.11) (x1.96) (x2.01) (x1.98) (x2.12) (x1.96) (x2.02) (x2.00) (x2.15) (x1.97) (x2.04) A 2.70 2.71 2.65 2.69 2.71 2.71 2.66 2.69 2.72 2.71 2.65 2.69 STAND (Ours) T 37.56 43.71 38.71 39.99 39.13 46.81 40.45 42.13 40.76 49.11 42.72 44.20 (x2.11) (x2.41) (x2.22) (x2.25) (x2.20) (x2.58) (x2.32) (x2.37) (x2.30) (x2.70) (x2.45) (x2.49) A 3.16 3.42 3.29 3.29 3.28 3.63 3.47 3.46 3.42 3.86 3.65 3.64
@@
In Figure˜1 and Table˜1, we evaluate STAND’s performance in multi-trajectory inference, where we generate multiple candidate answers by sequentially producing k independent reasoning traces and then aggregate the results.
@@
As shown in Figure˜1, STAND significantly improves decoding efficiency, achieving equivalent performance to plain decoding in less than 40% the time. Table˜1 provides detailed throughput and acceptance length comparisons across methods. STAND not only achieves the highest throughput but also maintains longer acceptance lengths compared to baselines. Importantly, both metrics improve as we increase the number of trajectories, making STAND’s speedup advantage more pronounced with increased compute scaling.
@@ 5.2 Evaluation on single-trajectory decoding
Table 2: Single-trajectory evaluations. We report the throughput (T) and acceptance length (A) for generating a single sequence with DeepSeek-R1-Distill-Qwen-7B and 14B. Best results are shown in bold.
AIME GPQA LCB Avg. DeepSeek-R1-Distill-Qwen-7B Plain T 26.63 31.34 27.75 28.57 Eagle-2 T 29.91 31.69 27.61 29.74 (x1.12) (x1.01) (x0.99) (x1.04) A 2.21 1.99 2.13 2.11 PLD T 44.34 42.84 43.40 43.53 (x1.67) (x1.37) (x1.56) (x1.52) A 1.72 1.64 1.59 1.65 ANPD T 46.18 54.05 44.79 48.34 (x1.73) (x1.72) (x1.61) (x1.69) A 1.88 1.82 1.80 1.83 SAM T 40.85 48.45 42.92 44.07 (x1.53) (x1.55) (x1.55) (x1.54) A 1.69 1.69 1.80 1.73 Recycle T 60.61 71.00 60.12 63.91 (x2.28) (x2.27) (x2.17) (x2.24) A 2.73 2.71 2.73 2.72 SAM+Recycle T 61.15 71.51 62.78 65.15 (x2.30) (x2.28) (x2.26) (x2.28) A 2.70 2.81 2.69 2.73 STAND (Ours) T 61.79 75.39 66.41 67.86 (x2.32) (x2.41) (x2.39) (x2.38) A 3.07 3.05 3.01 3.04 DeepSeek-R1-Distill-Qwen-14B Plain T 17.76 18.16 17.43 17.78 Eagle-2 T 25.38 24.86 21.89 24.04 (x1.43) (x1.37) (x1.26) (x1.35) A 2.72 2.44 2.51 2.56 PLD T 21.82 24.97 21.76 22.85 (x1.23) (x1.38) (x1.25) (x1.28) A 1.61 1.64 1.58 1.61 ANPD T 25.60 26.40 23.16 25.05 (x1.44) (x1.45) (x1.33) (x1.41) A 1.76 1.79 1.76 1.77 SAM T 23.26 25.38 22.36 23.67 (x1.31) (x1.40) (x1.28) (x1.33) A 1.63 1.65 1.63 1.64 Recycle T 33.71 38.91 33.85 35.49 (x1.90) (x2.14) (x1.94) (x2.00) A 2.77 2.73 2.71 2.74 SAM+Recycle T 34.35 37.53 34.45 35.44 (x1.93) (x2.07) (x1.98) (x1.99) A 2.67 2.72 2.70 2.70 STAND (Ours) T 34.52 38.71 34.86 36.03 (x1.94) (x2.13) (x2.00) (x2.03) A 2.91 3.00 2.93 2.95
While STAND is primarily designed to leverage information across multiple reasoning trajectories, we also evaluate its performance on single-trajectory generation, where the model only produces one long reasoning chain. As shown in Table˜2, STAND achieves both the highest acceptance length and throughput in most scenarios, demonstrating its effectiveness even when generating individual solutions.
5.3 Evaluation on diverse inference patterns
Table 3: Batch decoding evaluations. We report the throughput (T) and acceptance length (A) for batch decoding with DeepSeek-R1-Distill-Qwen-7B. Both speculative decoding approaches use trees optimized with OpenThoughts-114k. Best results are shown in bold.
AIME GPQA LCB Avg. Batch Size 4 Plain T 89.04 88.32 92.64 90.00 Recycle T 90.51 98.41 89.55 92.82 (x1.02) (x1.11) (x0.97) (x1.03) A 1.77 1.83 1.72 1.77 STAND (Ours) T 127.58 134.64 122.09 128.10 (x1.43) (x1.52) (x1.32) (x1.42) A 2.57 2.74 2.57 2.63 Batch Size 8 Plain T 111.62 106.23 114.73 110.86 Recycle T 99.18 105.69 99.2 101.36 (x0.89) (x0.99) (x0.86) (x0.91) A 1.66 1.75 1.63 1.68 STAND (Ours) T 148.08 154.68 149.41 150.72 (x1.33) (x1.46) (x1.30) (x1.36) A 2.70 2.86 2.69 2.75
Table 4: Test-time tree search evaluations. We report the throughput (T) and acceptance length (A) for performing Diverse Verifier Tree Search (DVTS) with DeepSeek-R1-Distill-Qwen-7B. Both speculative decoding approaches use trees optimized with OpenThoughts-114k. Best results are shown in bold.
AIME GPQA LCB Avg. Plain T 33.35 33.12 32.22 32.90 Recycle T 71.52 70.64 69.47 70.54 (x2.14) (x2.13) (x2.16) (x2.14) A 2.73 2.69 2.74 2.72 STAND (Ours) T 82.62 86.93 80.97 83.51 (x2.48) (x2.62) (x2.51) (x2.54) A 3.55 3.71 3.51 3.59
We extend our evaluation to more diverse and practical inference patterns, focusing on batch decoding and test-time tree search. In the batch decoding setup, multiple reasoning traces are generated simultaneously, and the N-gram drafters can only leverage the content produced up to the current decoding step. As shown in Table˜3, STAND consistently achieves significant speedups even with a batch size of 8, whereas Token Recycle provides only marginal improvements at batch size 4 and even degrades performance at batch size 8.
[…]
@@ 5.4 Ablations and analysis
Table 5: Effect of Stochastic Drafting. We report the throughput (T) and acceptance length (A) for generating 4 sequences with DeepSeek-R1-Distill-Qwen-7B.
AIME GPQA LCB Avg. Plain T 26.63 31.34 27.75 28.57 Deterministic T 62.13 73.67 63.44 66.41 (x2.33) (x2.35) (x2.29) (x2.32) A 2.94 2.98 2.90 2.94 Stochastic T 63.44 81.20 65.90 70.18 (x2.38) (x2.59) (x2.37) (x2.46) A 3.24 3.56 3.29 3.36 + Gumbel-Top-K T 64.99 83.47 69.70 72.72 (x2.44) (x2.66) (x2.51) (x2.55) A 3.21 3.48 3.30 3.33
@@
In Table˜5, we compare three drafting approaches: deterministic drafting, basic stochastic drafting (using PyTorch’s multinomial sampling), and our optimized stochastic drafting with Gumbel-Top-K. For fair comparison, we separately perform tree optimization for determinisic drafting and stochastic drafting. Stochastic drafting consistently achieves higher acceptance lengths across all tasks, resulting in improved throughput compared to deterministic drafting. Our Gumbel-Top-K optimization further improves performance by maintaining similar acceptance lengths while significantly reducing latency, leading to even higher throughput.
Table 6: Effect of tree optimization. Comparison of throughput and acceptance length when generating 4 sequences with DeepSeek-R1-Distill-Qwen-7B on two datasets: AIME-2024 and GPQA-Diamond. We compare two types of static trees: the heuristic trees from Token Recycle and our data-optimized trees, optimized on AIME-2024. Best results are shown in bold.
AIME GPQA (OOD) Heuristic Optimized Heuristic Optimized Throughput 59.96 64.99 77.32 83.47 Acc. Lens 3.17 3.21 3.35 3.48
@@
In Table˜6, we showcase the effectiveness of our tree optimization technique. We compare the performance of a heuristic tree originally used by Token Recycle (luo2024turning) with our tree, optimized on the AIME-2024 dataset. The results demonstrates that the optimized tree improves performance on both AIME-2024 and GPQA-Diamond, showcasing that the optimization not only works within the same dataset, but also generatlizes to out-of-domain (OOD) tasks.
@@
We analyze how different drafting approaches lead to different optimal tree structures by comparing trees optimized for STAND versus Token Recycle. As shown in Figure˜5, the tree optimized for STAND reaches greater depths, extending to 13 levels compared to 7 levels in the Token Recycle-optimized tree. This difference likely stems from STAND’s higher acceptance rate, which favors deeper, narrower tree structures under the same tree size budget.
@@ 6 Conclusion
In this work, we introduced STAND, a model-free speculative decoding approach that accelerates language model reasoning while maintaining accuracy. By utilizing reasoning trajectory redundancy and historical logit information, STAND significantly improves throughput over standard auto-regressive decoding and existing alternatives, offering an efficient solution for scaling AI reasoning systems.
@@
While STAND demonstrates strong performance across diverse inference patterns, all measurements were conducted using the HuggingFace implementation, which is less optimized than popular serving frameworks such as vLLM or SGLang. Although we expect the benefits of STAND to extend to these optimized frameworks, this has not yet been verified. In addition, the current N-gram lookup operation is implemented in Python, which may introduce a slowdown. Latency could be further reduced with a more optimized implementation of the N-gram module.
Acknowledgements
For WS and JS: This work was supported by Institute for Information & communications Technology Promotion(IITP) grant funded by the Korea government(MSIT) (No.RS-2019-II190075 Artificial Intelligence Graduate School Program (KAIST); No. RS-2024-00509279, Global AI Frontier Lab). This work was also supported by the NIPA(National IT Industry Promotion Agency), through the Ministry of Science and ICT (Hyperscale AI flagship project).
Appendix A Experimental Details
A.1 Initial tree for optimization